# Exploring and Evaluating Real-world CXL: Use Cases and System Adoption

### v1

@@ 5. CXL for Large Language Models
Figure 10. Data transfer bandwidth between the GPU (labeled as “g”) and CPU (labeled as “c”).
Bandwidth results. We measure memory access bandwidth by repeatedly and randomly copying data blocks between the GPU memory (labeled as “g”) and CPU memory (labeled as “c”) using various page allocation strategies. The size of the data blocks varies from 128 byte to 4GB, illustrated in Figure 10. We use membind (Andi Kleen ([n. d.]), SUSE Labs) to specify the memory device as the source or destination for data transfers.
@@
Counter-intuitively, using the CXL memory does not increase memory bandwidth for the GPU. As shown in Figure 10, the peak memory bandwidth is similar across various memory interleaving policies (the difference is less than 3%). This lack of bandwidth increase with the CXL memory is primarily due to the PCIe interconnect between the CPU and GPU acting as a performance bottleneck.
Figure 11. Data transfer latency between the GPU and CPU. The data size is 64 bytes.
Latency results. We develop a microbenchmark to measure the data transfer latency. The microbenchmark runs on CPU 1 (the CPU close to the CXL memory) and uses cudaMemcpy() for data transfer. The benchmark repeatedly transfers a 64-byte data (a cache block) between the CPU and GPU. The transfer happens 100K times, and we report average time for one transfer. We use membind, similar to the bandwidth tests. Figure 11 shows the results.
@@
The difference of data transfer latency between “GPU - CXL memory” and “GPU - CPU memory” is greater than the latency difference between “CPU - CXL memory” and “CPU - CPU memory”, shown in Figure 11. For example, accessing the CXL memory from the GPU is, on average, 500 $ns$ longer than accessing the CPU memory from the GPU. In contrast, accessing the CXL memory from the CPU is only 120 $ns$ longer than accessing the CPU memory from the CPU. The longer latency-difference from the GPU side comes from the longer data-path between the GPU and CXL memory.
@@
We study the implication of using the CXL memory for LLM using tensor offloading, as follows.
5.1. LLM Training
5.1.1. ZeRO-Offload Background
Figure 12. Overview of ZeRO-Offload in a training step.
@@ 5.1.1. ZeRO-Offload Background
ZeRO-Offload is a commonly employed in the industry to enable larger LLM training using smaller GPU memory. To save the GPU memory, ZeRO-Offload efficiently offloads full-precision model parameters, gradients, and optimizer states (e.g., momentum and variance) to the CPU memory, moving them back to GPU memory when needed. Figure 12 depicts the workflow of ZeRO-Offload. Specifically, it 1 performs forward and 2 backward computation on GPU; and 3 offloads gradients to the CPU memory during the backward step. To reduce the overhead of tensor movement between CPU and GPU 4, ZeRO-Offload performs optimization computation (e.g., the ADAM optimizer) on the CPU; and 5 moves updated parameters from the CPU memory to the GPU memory before the next forward step. This strategy minimizes data movement volume between the GPU and CPU memory for each training step. As a result, ZeRO-Offload enables 10$\times$ larger model training on a single GPU with 1.4$\times$ higher throughput.
@@
Figure 13. Performance with various interleaving policies and model sizes for BERT and GPT2.
5.1.2. Using CXL Memory
@@
We evaluate two LLMs: BERT (Devlin et al., 2018) and GPT2  (Radford et al., 2019). For BERT, we consider three configurations: 110 million (base), 340 million (medium), and 4 billion (large) parameters. For GPT2, we evaluate models with 4, 6, and 8 billion parameters. Figures 13 shows the performance with various interleaving policies and model sizes. We use the notation “bs=effective batch size@model size” to represent the batch size and number of model parameters. For a given model size, the batch size is chosen to be the maximum without causing an out-of-memory (OOM) error on the GPU. To better understand the performance, we break it down to the “optimization step” (i.e., the ADAM optimizer on the CPU, which is exposed to the critical path) and “data movement” (i.e., the gradient transfer from the GPU to the CPU and the parameter transfer from the CPU to the GPU). The memory capacities for various interleaving policies (i.e., LDRAM only, LDRAM+CXL, LDRAM+RDRAM, and “interleave all”) are 196GB, 324GB, 392GB, and 520GB. Figures 14 shows the results, including the data movement exposed to the critical path.
@@
Figure 14. Performance of the optimizer and data movement in BERT and GPT2 under various interleaving policies. The percentage numbers are in terms of total training time.
@@
Evaluation results. LLM training observation 1. Using the CXL memory bring little performance improvement or even negative performance impact to ZeRO-Offloading.
@@
Take GPT2 as an example. Figure 13 shows that the performance difference between LDRAM+RDRAM, LDRAM+CXL, and “interleave all” is less than 5% for the 4B and 6B models. For the 8B model, LDRAM outperforms ‘interleave all” by 14%, and LDRAM+RDRAM outperforms LDRAM+CXL by 16%. There is no performance benefit of using CXL.
@@
Figure 14 reveals three reasons for the above observation. (1) The data movement takes a rather small portion of the training time in GPT2 (less than 5% in all cases). (2) Because of the bottleneck in the extra PCIe, the data movement does not get benefits from the better bandwidth offered by interleaving CXL and DDR (the CPU memory). Instead, the data movement suffers from longer access latency of the CXL memory. As a result, compared with using LDRAM, using CXL actually increases data transfer time and has a minor impact on the training time. (3) The optimizer takes a larger portion of the training time, compared to the data movement. The optimizer happens on the CPU and is sensitive to memory latency. Using CXL increases its execution time (2%-18%), compared to using LDRAM. When the batch size is small, he optimizer takes a significant portion of the training time (e.g., 31% when bs=3@8B). In such cases, optimizer slowdown substantially decreases overall training throughput. For example, when bs=3@8B, using “interleave all” performs worse than LDRAM+RDRAM by 11% due to the worse performance in the optimizer by 8%. We have the similar observations for BERT.
@@
We expect that using a larger model (such as GPT-3 with 175 billion parameters) would result in data movement taking a larger portion of the training time. Unfortunately, we cannot evaluate such a large model because of the limited memory capacity in our GPU. In addition, after reducing the data path between the GPU and CXL memory, the CXL memory can play a bigger role in reducing training time. Also, using the interleaving policy for the optimizer is not good for performance because of the latency sensitive nature of the optimizer. Using the first touch or “preferred policy”, the optimizer performance can be improved.
5.2. LLM Inference
5.2.1. FlexGen Background
@@
LLM inference is memory-consuming. FlexGen (Sheng et al., 2023) is a cutting-edge framework designed for LLM inference with constrained GPU memory capacity. To address the memory capacity limitation, FlexGen offloads model parameters, KV cache, and activations to the host CPU memory hierarchy (including DRAM and NVMe SSD). Figure 15 shows the workflow of FlexGen. The LLM inference consists of two stages: prefill and decode.
@@
Figure 15. Overview of LLM inference with FlexGen.
@@
5.2.2. Using CXL Memory
Evaluation setup. We evaluate LLaMA (Touvron et al., 2023) with 65 billion parameters and OPT (Zhang et al., 2022) with 66 billion parameters. The lengths of the input prompts and output tokens are standardized at 2,048 and 256, respectively. The evaluation platform has 196 GB LDRAM, 196 GB RDRAM, 128 GB CXL, and 128 GB NVMe. Leveraging GRUB mmap and numactl (Andi Kleen ([n. d.]), SUSE Labs), we build the CPU memory hierarchy with various capacities and media types.
Figure 16. Comparison LLM inference throughput across memory systems, each with 324 GB capacity.
@@ 5.2.2. Using CXL Memory
Table 3. LLM inference configuration for evaluation. “BS” and “c” stands for “batch size” and “KV cache”, respectively. According to offloading policy in FlexGen, all weights and activations are stored in CPU memory.
@@
Figure 17. Comparison of LLM inference throughput across various memory systems with different capacities.
@@
Figure 17 illustrates the inference throughput with various capacities of the CPU memory. Table 3 summarizes the offloading policy search results with those capacities. The LLM batch size scales up with the increased memory system capacity. Specifically, with LDRAM + CXL, LDRAM + RDRAM, and LDRAM + RDRAM + CXL, the batch size increases by 1.14$\times$, 1.85$\times$, and 3$\times$ for LLaMA, and 2.11$\times$, 3.44$\times$ and 6.11$\times$ for OPT, respectively, compared to the case of utilizing LDRAM only. The overall throughput increases by 28%, 81% and 86% on average respectively, compared to LDRAM only.
@@
6. Memory Tiering based on Page Migration
@@
Memory tiering is an application-transparent solution to integrate the CXL memory into the existing memory systems. Treating the CXL memory as a memory tier, existing memory tiering solutions (Corbet, [n. d.]; Verma., 2022; Ren et al., 2024; Maruf et al., 2023; Raybuck et al., 2021; Lee et al., 2023) rely on memory profiling to count memory accesses at the page level. Then, those solutions move frequently accessed (hot) pages to the fast memory tier, and demote less frequently access (cold) pages to the slow memory tier. In contrast to the interleaving studied in Sections 4-5, which are static page placement solutions, the memory tiering solutions are dynamic.
@@
We use the system A for evaluation. We limit the capacity of the fast memory tier (i.e., LDRAM) while the slow memory tier (i.e., CXL) has unlimited capacity. We evaluate three state-of-the-art memory tiering solutions as follows. We also evaluate No Balance (between NUMA nodes) representing the static page placement without migration. Our study is featured with analyzing the interplay between page migration and interleaving.
  • AutoNUMA (Corbet, [n. d.]) tracks memory accesses by unmapping pages to generate hint faults, and periodically scans the application address space to find hint faults. AutoNUMA promotes pages using a NUMA-distance-based, static memory-access threshold to avoid unnecessary page migration. AutoNUMA is enabled by setting $numa\_balancing$ to 1 in /proc/sys/kernel (the default Linux setting).
@@
  • Tiering-0.8 (Verma., 2022) also uses hint faults to identify hot pages. Unlike AutoNUMA, Tiering-0.8 dynamically adjusts the promotion criteria to throttle migration traffic. It considers the recency of page accesses to decide page promotion and demotion. Tiering-0.8 is enabled by setting $numa\_balancing$ to 2, and it uses the Linux default promotion threshold for initialization.
  • TPP (Maruf et al., 2023) utilizes both hint faults and the active list in the OS to determine page migration, considering both access recency and page activeness for promotion.
@@
Figure 18. Execution time with various page migration and static page placement solutions for memory-intensive applications.
Table 4. NUMA statistics with various page migration and static page placement solutions.
AutoNUMA Tiering-0.8 TPP
hint_faults migrated_pages hint_faults migrated_pages hint_faults migrated_pages
BTree First touch 336,469,816 3 125,802,092 46 4,654,013,085 5
Interleave 6,092 492 6,916 51 5,092 1
PageRank First touch 31,793,182 117 63,893,091 7,653,699 5,894,907,183 6,929,247
Interleave 28,398 10,050 613,906 4,412 165,721 15
Graph500 First touch 66,495,782 118,621 398,259,304 6,134 36,393,830,981 190
Interleave 4,381,521 124 282,087 6,114 74,432,220 1,569
Silo First touch 146,665,022 21,494 220,094,592 3,264 1,588,732,736 1,497,420
Interleave 11,428 10 5,581,434 14,512 60,638 0
@@
Figure 19. Execution time with various page management solutions for HPC applications.
6.1. Page Migration with Application-Level Interleaving

### v2

@@ IV CXL for Large Language Models
Figure 5: Data transfer bandwidth between GPU (i.e., “g”) and CPU (i.e., “c”).
Bandwidth results. We measure memory access bandwidth by repeatedly and randomly copying data blocks between the GPU memory (labeled as “g”) and CPU memory (labeled as “c”) using various page allocation strategies. The size of the data blocks varies from 128 byte to 4GB, illustrated in Figure 5. We use membind [5] to specify the memory device as the source or destination for data transfers.
@@
Counter-intuitively, using the CXL memory does not increase the memory bandwidth for the GPU. As shown in Figure 5, the peak memory bandwidth is similar across various memory interleaving policies (the difference is less than 3%). This lack of bandwidth increase with the CXL memory is primarily due to the PCIe interconnect between the CPU and GPU acting as a performance bottleneck.
Figure 6: 64-byte data transfer latency between the GPU and CPU.
Latency results. We develop a microbenchmark to measure the data transfer latency. The microbenchmark runs on CPU 1 (the CPU close to the CXL memory) and uses cudaMemcpy() for data transfer. The benchmark repeatedly transfers a 64-byte data (a cache block) between the CPU and GPU. The transfer happens 100K times, and we report the average time for one transfer. We use membind, similar to the bandwidth tests. Figure 6 shows the results.
@@
The difference of data transfer latency between “GPU - CXL memory” and “GPU - CPU memory” (in Figure 6) is greater than the latency difference between “CPU - CXL memory” and “CPU - CPU memory” (in Figure 2). For example, accessing the CXL memory from the GPU is, on average, 500 $ns$ longer than accessing the CPU memory from the GPU. In contrast, accessing the CXL memory from the CPU is only 120 $ns$ longer than accessing the CPU memory from the CPU. The longer latency-difference from the GPU side comes from the longer data-path between the GPU and CXL memory.
@@
We study the implication of using the CXL memory for LLM training and inference using tensor offloading on System A, as follows.
IV-A LLM Training
IV-A1 ZeRO-Offload Background
Figure 7: Overview of ZeRO-Offload in a training step.
@@ IV-A1 ZeRO-Offload Background
ZeRO-Offload [54] is commonly employed in the industry to enable larger LLM training using smaller GPU memory. To save the GPU memory, ZeRO-Offload efficiently offloads full-precision model parameters, gradients, and optimizer states (e.g., momentum and variance) to the CPU memory, moving them back to GPU memory when needed. Figure 7 depicts the workflow of ZeRO-Offload. Specifically, it 1 performs forward and 2 backward computation on GPU; and 3 offloads gradients to the CPU memory during the backward step. To reduce the overhead of tensor movement between CPU and GPU 4, ZeRO-Offload performs optimization computation (e.g., the ADAM optimizer) on the CPU; and 5 moves updated parameters from the CPU memory to the GPU memory before the next forward step. This strategy minimizes data movement volume between the GPU and CPU memory for each training step. As a result, ZeRO-Offload enables 10$\times$ larger model training on a single GPU with 1.4$\times$ higher throughput.
@@
Figure 8: Performance with various interleaving policies and model sizes for BERT and GPT2.
IV-A2 Using CXL Memory
@@
We evaluate two LLMs: BERT [14] and GPT2  [49]. For BERT, we consider three configurations: 110 million (base), 340 million (medium), and 4 billion (large) parameters. For GPT2, we evaluate models with 4, 6, and 8 billion parameters. Figure 8 shows the performance with various interleaving policies and model sizes. We use the notation “bs=effective batch size@model size” to represent the batch size and number of model parameters. For a given model size, the batch size is chosen to be the maximum without causing an out-of-memory (OOM) error on the GPU. To better understand the performance, we break it down to the “optimization step” (i.e., the ADAM optimizer on the CPU, which is exposed to the critical path) and “data movement” (i.e., the gradient transfer from the GPU to the CPU and the parameter transfer from the CPU to the GPU). The memory capacities for various interleaving policies (i.e., LDRAM only, LDRAM+CXL, LDRAM+RDRAM, and “interleave all”) are 196GB, 324GB, 392GB, and 520GB. Figure 9 shows the results, including data movement exposed to the critical path.
@@
Figure 9: Performance of the optimizer and data movement under various interleaving. The percentage numbers are in terms of total training time.
@@
Evaluation results. LLM training observation 1. Using the CXL memory brings little performance improvement or even negative performance impact to ZeRO-Offloading.
@@
See GPT2. Figure 8 shows that the performance difference between LDRAM+RDRAM, LDRAM+CXL, and “interleave all” is less than 5% for the 4B and 6B models. For the 8B, LDRAM outperforms ‘interleave all” by 14%, and LDRAM+RDRAM outperforms LDRAM+CXL by 16%. There is no performance benefit of using CXL.
@@
Figure 9 reveals three reasons for the above observation. (1) The data movement takes a rather small portion of the training time in GPT2 (less than 5% in all cases). (2) Because of the bottleneck in the extra PCIe, the data movement does not get benefits from the better bandwidth offered by interleaving CXL and DDR (the CPU memory). Instead, the data movement suffers from longer access latency of the CXL memory. As a result, compared with using LDRAM, using CXL actually increases data transfer time and has a minor impact on the training time. (3) The optimizer takes a larger portion of the training time, compared to the data movement. The optimizer happens on the CPU and is sensitive to memory latency. Using CXL increases its execution time (2%-18%), compared to using LDRAM. When the batch size is small, the optimizer takes a significant portion of the training time (e.g., 31% when bs=3@8B). In such cases, optimizer slowdown substantially decreases overall training throughput. For example, when bs=3@8B, using “interleave all” performs worse than LDRAM+RDRAM by 11% due to the worse performance in the optimizer by 8%. We have the similar observations for BERT.
@@
We expect that using a larger model (such as GPT-3 with 175 billion parameters), which has larger numbers of parameters and gradients, would result in data movement taking a larger portion of training time to transfer parameters and gradients. Unfortunately, we cannot evaluate such a large model due to the limited memory capacity on our GPU. In addition, after reducing the data path between the GPU and CXL memory, the CXL memory can play a bigger role in reducing training time. Also, using the interleaving policy for the optimizer is not good for performance because of the latency sensitive nature of the optimizer. Using the first touch or “preferred policy”, the optimizer performance can be improved.
IV-B LLM Inference
IV-B1 FlexGen Background
@@
LLM inference is memory-consuming. FlexGen [63] is a cutting-edge framework designed for LLM inference with constrained GPU memory capacity. To address the memory capacity limitation, FlexGen offloads model parameters, KV cache, and activations to the host CPU memory hierarchy (including DRAM and NVMe SSD). Figure 10 shows the workflow of FlexGen. The LLM inference consists of two stages: prefill and decode.
@@
Figure 10: Overview of LLM inference with FlexGen.
@@
IV-B2 Using CXL Memory
Evaluation setup. We evaluate LLaMA [69] with 65 billion parameters and OPT [93] with 66 billion parameters. The batch sizes for evaluation are determined by performing a linear programming-based [66] policy search to maximize inference throughput while respecting memory and hardware constraints. The lengths of the input prompts and output tokens are standardized at 2,048 and 256, respectively. The evaluation platform has 196 GB LDRAM, 196 GB RDRAM, 128 GB CXL, and 128 GB NVMe. Leveraging GRUB mmap and numactl [5], we build the CPU memory hierarchy with various capacities and media types.
Figure 11: Comparison LLM inference throughput across memory systems, each with 324 GB capacity.
@@ IV-B2 Using CXL Memory
TABLE II: LLM inference configuration for evaluation. “BS” and “c” stands for “batch size” and “KV cache”, respectively. According to offloading policy in FlexGen, all weights and activations are stored in CPU memory.
@@
Figure 12: Comparison of LLM inference throughput across various memory systems with different capacities.
@@
Figure 12 illustrates the inference throughput with various capacities of the CPU memory. Table II summarizes the offloading policy search results with those capacities. The LLM batch size scales up with the increased memory system capacity. Specifically, with LDRAM + CXL, LDRAM + RDRAM, and LDRAM + RDRAM + CXL, the batch size increases by 1.14$\times$, 1.85$\times$, and 3$\times$ for LLaMA, and 2.11$\times$, 3.44$\times$ and 6.11$\times$ for OPT, respectively, compared to the case of utilizing LDRAM only. The overall throughput increases by 28%, 81% and 86% on average respectively, compared to LDRAM only.
@@
V HPC Workloads
We analyze HPC workload performance, focusing on seven workloads from NPB [15] and XSBench benchmark [70], which require processing a large working memory set using multiple threads, are summarized in Table III. They collectively cover the most common and representative HPC applications [6]. We use various memory allocation policies, including preferred and interleaving. The preferred policy for a specific memory node indicates that the memory is allocated in that memory node first; when that memory node runs out of space, the page allocation goes to another memory node closet to the CPU according to the NUMA distance. The interleaving policy indicates that pages are allocated among memory nodes in a round robin fashion. We present evaluation results on the system A, because we have limited accesses to B and C.
TABLE III: HPC workloads for evaluation. “BW” stands for “bandwidth”.
BT Dense linear algebra Unit-strided memory accesses from dense matrices Class E 166 GB u(39.6G), rsh(39.6G), forcing(39.6G)
LU Sparse linear algebra Indexed loads and stores from compressed matrices Class E 134 GB u(39.6G), rsd(39.6G)
CG Sparse linear algebra Irregular memory accesses based on indirect indexing Class E 134 GB a(48.9G)
MG Structured grids Dynamic updates based on subdivided regular grids Class E 210 GB v(64.2G), r(73.4G)
SP Structured grids Intense floating-point computations for linear equations Class E 174 GB u(39.6G), rsh(39.6G), forcing(39.6G)
FT Spectral method Bandwidth-consuming matrix transpose Class D 80 GB u0(32.0G), u1(32.0G)
V-A Evaluation Results
We have the following observations unseen in the existing CXL evaluation [67, 68].
HPC observation 1. When the interleaving involves the CXL memory, we can save LDRAM by using RDRAM.
Figure 13 shows the results with various interleaving policies, and the benchmarks run on CPU 0. We can see that the performance difference between the interleaving RDRAM+CXL and interleaving LDRAM+CXL is less than $9.2\%$ for all benchmark. In general, we can save LDRAM while achieving similar performance by interleaving RDRAM+CXL instead of interleaving LDRAM+CXL.
The above results are because of large performance gap between the CXL memory and DDR (e.g., 2.1×2.1$\times$2.1 × and 1.2×1.2$\times$1.2 × longer in memory access latency, compared to LDRAM and RDRAM respectively.): because of data dependency and limited hardware resources (e.g., Miss Status Handling Registers (MSHR) and the queues in MC), the performance is highly impacted by the slow CXL memory and irrelevant whether the LDRAM or RDRAM is utilized.
Figure 13: The performance of various interleaving policies for HPC applications.
HPC observation 2. Bandwidth-sensitive and latency-sensi
Figure 14 shows the results for MG (bandwidth-sensitive) and CG (latency-sensitive). For MG, we see that when the number of threads increases from 4 to 32, the performance of “interleave all” (i.e., interleaving between LDRAM, RDRAM and CXL, which achieves the highest bandwidth) is consistently better than that of CXL preferred by 10%-85%. Also, for CG, the performance of “interleave all” performs worse than that of CXL preferred by up to 1.6$\times$, because using CXL preferred, consecutive memory accesses tend to fall into the same memory node, which is favored by CG for high performance (see more discussions for HPC observation 3).
Figure 14: The scalability study for CG and MG. The performance is normalized by that of using LDRAM only.
HPC observation 3. The CXL memory can show unexpected high performance for the latency-sensitive applications with random memory accesses.
The CG results in Figures 14 and  13 support the above observation. CG is a latency-sensitive application because of its indirect indexing-based memory accesses. In Figure 14, CXL preferred performs better than using RDRAM-only by $10.9\%$-$57.2\%$ when the number of threads is 4-20. This seems to be counter-intuitive, because the CXL memory has higher latency than RDRAM. This superior performance comes from the optimization in the CXL device or customized caching policy in the processor for expensive, CPU-less memory accesses [67], and this optimization is especially effective for CG-style workloads. When the number of threads is larger than 20, the CXL memory accesses become more intensive, and CXL inferior performance becomes more obvious.
V-B Object-Level Interleaving
Instead of generally interleaving pages across the entire application (named uniform interleaving), we propose a method to interleave pages at the data object level. This enables fine-grained control over how pages are interleaved. Since different data objects have different access patterns, using the fine-grained control allows the bandwidth-sensitive object to be accessed with high bandwidth, while the latency-sensitive object is allocated locally for high performance.
Interleaving method. To implement object-level interleaving, we employ numa_alloc_interleaved_subset() [5] in Linux, which allows for the allocation of a data object in an interleaved manner among specific NUMA nodes. We use two criteria to select data objects to use the interleaving policy.
With high thread-level parallelism, memory accesses to the above data objects are more sensitive to memory bandwidth. Besides the above data objects, the memory allocation for other objects uses the “preferred” policy. The last column in Table III summarizes those bandwidth-hungry data objects selected for interleaving.
We evaluate the effectiveness of our object-level interleaving using the following approach. In each test, we run the workload on CPU 0 using both LDRAM (memory node 0) and CXL memory. LDRAM is limited to either 64GB or 128GB by using GRUB mmap, and the memory consumption of all applications exceeding 64GB, which allows us to assess the cases with both sufficient and insufficient LDRAM. The CXL memory is consistently 128GB. We make the following object-level interleaving observations (abbreviated as OLI observations).
OLI observation 1. When the local DRAM is sufficient (with 128 GB), the object-level interleaving performance is similar to that with LDRAM preferred, and consistently outperforms the uniform interleaving in all scenarios.
Figure 15(a) shows that OLI performance is 65% better than the uniform interleaving on average, demonstrating the effectiveness of OLI to meet the diverse requirements of latency-sensitive and bandwidth-sensitive objects. Additionally, we observe that by using OLI, we can effectively reduce the fast memory size by an average of 32% (and up to 40% for FT), while still achieving similar performance as LDRAM preferred. The performance difference between OLI and LDRAM preferred is less than 1% on average for all HPC workloads, except XSBench, when LDRAM is sufficient.
Figure 15: (a) Performance speedup of various memory allocation strategies compared to “LDRAM preferred”, using sufficient LDRAM (128GB). (b) Performance speedup of various memory allocation strategies, compared to “LDRAM preferred”, using insufficient LDRAM (64GB).
Figure 15(b) shows that on average, OLI performs 1.42$\times$ better than LDRAM preferred (and up to 2.35$\times$ for BT). It also outperforms the uniform interleaving by 1.32$\times$ on average (and up to 1.84$\times$ for LU). See the following reasons. (1) The performance of LDRAM preferred is highly related to when the data objects are allocated. If LDRAM is full, latency-sensitive objects might end up in the slower CXL memory. (2) The bandwidth-sensitive data objects do not have opportunities to be allocated on both LDRAM and CXL for high bandwidth.
In XSBench, LDRAM preferred performs better than both the uniform and object-level interleaving. This is because most of the memory accesses in XSBench are concentrated in a small, latency-sensitive memory set.
Takeaway: The CXL memory offers additional bandwidth, but HPC workloads may not benefit from it directly. The uniform interleaving can undermine performance, while the data object-level interleaving delivers performance comparable to or better than the default LDRAM-centric page allocation while saving fast memory size.
VI Memory Tiering based on Page Migration
@@
Memory tiering is an application-transparent solution to integrate the CXL memory into the existing memory systems. Treating the CXL memory as a memory tier, existing memory tiering solutions [13, 73, 55, 44, 50, 32] rely on memory profiling to count memory accesses at the page level. Then, those solutions move frequently accessed (hot) pages to the fast memory tier, and demote less frequently access (cold) pages to the slow memory tier. In contrast to the interleaving studied in Sections IVV, which are static page placement solutions, the memory tiering solutions are dynamic.
We use the system A. We limit the capacity of the fast memory (i.e., LDRAM), while the capacity of the slow memory (i.e., CXL) remains the same.
@@
We evaluate three state-of-the-art memory tiering solutions as follows. We also evaluate No Balance (between NUMA nodes) representing the static page placement without migration. Our study is featured with analyzing the interplay between page migration and interleaving.
  • AutoNUMA [13] is the default NUMA-balancing policy in Linux. AutoNUMA counts memory accesses by manipulating an access bit in PTE. When a page is accessed, a NUMA hint fault is triggered. By tracking these faults, AutoNUMA determines the origin of memory accesses and relocates pages to minimize their distance from the computing processes. AutoNUMA is enabled by setting numa_balancing to 1 in /proc/sys/kernel.
@@
  • Tiering-0.8 [73] is a recent Linux patch which uses the similar hint fault-based profiling as AutoNUMA. Unlike AutoNUMA, it considers the recency of page accesses based on the re-fault interval to identify hot pages. Additionally, Tiering-0.8 dynamically adjusts the page promotion criteria to throttle migration traffic and save memory bandwidth. Tiering-0.8 is enabled by setting numa_balancing to 2.
  • TPP [44] uses hint faults to determine page migration. Upon encountering hint faults, TPP decides to promote a page based on its presence on the LRU list in Linux.
@@
Figure 16: Execution time with various page migration and static page placement solutions for memory-intensive applications.
@@
Figure 17: Execution time with various page management solutions for HPC applications.
TABLE IV: Comparison of Performance Studies on Genuine CXL
Paper HPC Study LLM Study Memory Tiering Performance Optimization
Training Inference Solutions Evaluated Identify Limitations
MICRO23 [67] No N/A N/A TPP N/A BW-aware page allocation
EuroSys24 [68] No N/A CPU-based AutoNUMA N/A N/A
SupMario [41] Yes N/A CPU-based TPP Yes Predictive interleaving & regulated migration
This work Yes GPU-based GPU-based TPP, AutoNUMA, Tiering-0.8 Yes Object-level interleaving
VI-A Page Migration with Uniform interleaving